OpenStudy (anonymous):

What is the approximate surface area of a right hexagonal prism with a base perimeter of 96 meters and a height of 10 meters? (Use S = ap + ph, whe4re a is the apothem of the base, p is the perimeter of the base, and h is the height of the prism.) Show all work.

OpenStudy (anonymous):

*where

OpenStudy (anonymous):

Wow, I come back to see no one has helped you

OpenStudy (anonymous):

96/6=16 so each line on the base is 16 for vertical walls 16x10= area of one wall, hexagon mean six walls so area of six walls is 960 Area of hexagon =(3)^1.5/2(side)^2 (36)3^(1.5)/2= area of base so base x2 is both top and bottom = (36)3^1.5 (36)3^1.5+960= surface area

OpenStudy (anonymous):

I get 2290 meters squared

OpenStudy (anonymous):

Your calculations seem strange

OpenStudy (anonymous):

scsilver, you found the apothem and used the given formula to solve it, right?